Find the area on the surface of a circular pond that has a diameter of 12m .

2] The pond is surrounded by path of width 2m if it costs 55dollar per meter square to pare the path with tiles find the cost.

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

12 m diameter
6 m radius
A=pi*r^2=36 pi m^2
circumference of pond is pi*D=12 pi meters
The circumference of the path: The path is a circle with diameter 16 m, 2 on each side of the pond, so radius is 8 m
The area of pond+path is 64 pi m^2. The area of the pond alone is 36 pi m^2, so the path is 28 pi m^2=87.96 m^2 or 88 m^2
The cost will be 88m^2*$55/m^2=$4840. My eyes saw a $ in place of another 5. Sorry!

Answer by ikleyn(52784)

.
Find the area on the surface of a circular pond that has a diameter of 12m .

2] The pond is surrounded by path of width 2m
if it costs 55dollar per meter square to pare the path with tiles find the cost.
~~~~~~~~~~~~~~~~~


The correct answer is THIS:

    At the price of 55 dollars per square meter,

    the cost of paving is  88*55 = 4840 dollars.